Reports 1-1 of 1 Clear search Modify search
AdV-COM (1/√f noise)
mwas - 20:39 Friday 23 January 2026 (68560) Print this report
EDB OMC lock on HOM and optical spring SNR increase test

Figure 1 summarizes the realignment of the EDB OMC using the USB TEM00, and then the lock on the carrier HOM between order 2 and 9

14:35 UTC Locked on EDB OMC 56MHz USB TEM00. Power around 0.5mW on B1t.
Realignign the OMC
15:06 UTC (6min) lock on USB TEM00, between 0.55mW and 0.6mW. With low
jitter peaks at a few hundred Hz

15:22 UTC (15min) lock on carrier order 2 mode, unstable, jumping back - Figure 2
and forth between the vertical and horizontal mode, the later having
double the power

15:45 UTC (10min) lock on carrier order 3 mode, horizontal peak - Figure 3

16:03 UTC (10min) lock on carrier order 4 mode, horizontal peak - Figure 4

16:22 UTC (10min) lock on carrier order 5 mode, vertical peak

16:37 UTC (10min) lock on carrier order 5 mode, horizontal peak that has 3 times more power

16:58 UTC (10min) lock on carrier order 6 mode, horizontal peak - Figure 5

17:18 UTC (10min) lock on carrier order 7 mode, horizontal peak

17:48 UTC (10min) lock on carrier order 8 mode, horizontal peak - Figure 6

18:08 UTC (20min) lock on carrier order 9 mode, horizontal peak - Figure 7

Restarting the EDB OMC scan with  freq 0.0005Hz, ampl 0.5, offset 23.27

19:00 UTC increased DARM line amplitude (74.4Hz) by a factor 2 and reduce OS calibration by a factor 2, using a Cm command.

Figure 8. The change in DARM line clearly broke the LSC_DCP_moni_mad_cal monitor of the DCP frequency. The Hrec is still working. Note that the Hrec measurement of the DCP follows almost exactly the SRCL_INPUT signal, after the factor 2 increase in DARM line amplitude it appears more stable. To be confirmed with a few more hours of data if the DCP really becomes more stable with a higher SNR for the optical spring signal that control SRCL set point.

Images attached to this report
Comments to this report:
mwas - 8:10 Saturday 24 January 2026 (68561) Print this report

Figure 1. Increasing the DARM line by a factor 2 has been effective at making the optical response pole (DCP) more stable. It also reduced the SRCL INPUT fluctuations. 

Figure 2. The DCP frequency and SRCL INPUT after the change (purple) are less correlated and have a lower spectrum below 1mHz than the previous night (blue). So there is improvement, but SRCL INPUT still dominates the slow fluctuations of the DCP with a coherence of ~60%. 

Having a larger DARM line create a large bump at 74Hz in the sensitivity. This is acceptable during the current commissioning period, but wouldn't be acceptable during a science run. In the long term a DARM line at lower frequency should provide a better signal for SRCL, as the radiation pressure effects scale as 1/f^2, so should be much larger and clearly visible in B7 for a lower frequency DARM line.

Images attached to this comment
ruggi - 12:45 Saturday 24 January 2026 (68562) Print this report

I think not only the DCP poor stability matters, but also the fact that we are not working in the maximum. The attached figure shows  SRCL vs DCP correlation and the presence of an offset. The way we compute SRCL set is an indirect estimation of optical spring, assuming that the power in the arms is correlate to DARM line only by radiation pressure, but any other coupling mechanism could create an offset.

I think we could try to demodulate the DCP wrt a line on SRCL and use it to define the optimal SRC longitudinal working, as we do for the angular working point. The line could stay below 10 Hz as for the angular lines.

Images attached to this comment
Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×